# appendix,by

### v1

@@ 1.2 Tropical geometry
In what follows, we will use results from tropical geometry, for which we refer the reader to [MS15]. In the cited source, all value groups are Archimedean, which is typically not the case in this paper. Nevertheless, all the results we will use from [MS15] still hold in our setting; since this sometimes requires justification, and because it may be of independent interest, we gather in this subsection the results that we will need. We especially want to highlight that many of these are, or can be reduced to, first-order statements; therefore, by the model-theoretic properties of $\mathsf{ACVF}$, they do not depend on the rank of the valuation.
@@
    The tropicalisation [^2] This definition does not require the existence of a cross-section. $\operatorname{trop}(f)\colon\Gamma^{n}\to\Gamma$ is defined as
@@
    The initial form of $f$ with respect to $\gamma$ is the Laurent polynomial over the residue field defined as $\operatorname{in}_{\gamma}(f)\coloneqq\sum_{\mathbf{u}\in S^{\prime}}\operatorname{res}(c_{\mathbf{u}}s(-v(c_{\mathbf{u}})))x^{\mathbf{u}}$, where $S^{\prime}\coloneqq\{\mathbf{u}\in S\mid\forall\mathbf{u}^{\prime}\in S,v(c_{\mathbf{u}})+\langle\mathbf{u},\gamma\rangle\leq v(c_{\mathbf{u}^{\prime}})+\langle\mathbf{u}^{\prime},\gamma\rangle\}$.
@@
    If $V\subseteq\mathbb{G}_{\mathrm{m}}^{n}$ is an algebraic subvariety defined over $K$ and $I$ is the ideal of Laurent polynomials over $K$ vanishing on $V$, then we denote by $\operatorname{in}_{\gamma}(V)$ the algebraic subvariety of $\mathbb{G}_{\mathrm{m}}^{n}$ (defined over $k$) associated to $\operatorname{in}_{\gamma}(I)$. We call this the initial variety of $V$ with respect to $\gamma$.
Note that initial forms depend on the choice of a cross-section, even if this is suppressed from the notation. Below, if we mention initial forms without specifying a cross-section, we assume implicitly that a one exists and has been fixed. Recall that cross-sections always exist if the value group is divisible, e.g. in the algebraically closed case, see for example [ADH17, Lemma 3.3.32].
@@
    In order to define initial forms one does not really need the cross-section $s$, rather the corresponding angular component $\operatorname{res}(xs(-v(x)))$. We mostly use the cross-section as that is customary in the literature, but observe that many results in tropical geometry (say from [MS15, Sections 2 and 3]) generalise to initial forms when they are defined using an arbitrary angular component.
@@
It is almost immediate to see that the codimension 1 case of the Fundamental Theorem of Tropical Geometry, also known as Kapranov’s Theorem [MS15, Theorem 3.1.3], is a first-order statement. In higher codimension we will need a bound of Joswig and Schröter on the minimum degree of a tropical basis.
@@
Let $K$ be a valued field and $I$ an ideal in $K[x^{\pm 1}]$. A tropical basis of $I$ is a finite generating subset $\mathcal{T}\subseteq I$ such that for every $\gamma\in\Gamma^{n}$, if there is $f\in I$ such that $\gamma$ is not a tropical root of $\operatorname{trop}(f)$, then there is $g\in\mathcal{T}$ such that $\gamma$ is not a tropical root of $\operatorname{trop}(g)$.
@@
Let $K$ be an algebraically closed valued field with value group $\mathbb{Q}$. For all finite $S\subseteq\mathbb{Z}^{n}$ there is a finite $H\subseteq\mathbb{N}^{n}$ such that the following holds. Let $I\subseteq K[x^{\pm 1}]$ be an ideal and $\mathcal{G}$ a finite set generating $I$ and such that every polynomial in $\mathcal{G}$ can be written as $\sum_{\mathbf{u}\in S}c_{\mathbf{u}}x^{\mathbf{u}}$ (with possibly some $c_{\mathbf{u}}=0$). Then there is a tropical basis $\mathcal{T}\supseteq\mathcal{G}$ of $I$ such that every polynomial in $\mathcal{T}$ can be written as $\sum_{\mathbf{u}\in H}b_{\mathbf{u}}x^{\mathbf{u}}$.
@@
Given the ideal $I$, denote by $I_{\mathrm{proj}}$ its homogenisation, that is, the homogeneous ideal in $K[x_{0},\ldots,x_{n}]$ obtained by homogenising the elements of $K[x_{1},\ldots,x_{n}]\cap I$.
Because $\Gamma(K)=\mathbb{Q}$, every ideal of $K[x_{0},\ldots,x_{n}]$ is defined over a subfield of $K$ with discrete valuation. Hence we may apply [JS18, Theorem 10] to $I_{\mathrm{proj}}$ and obtain that there are $d_{0}$ —depending only on $n$ and on the degree of $I_{\mathrm{proj}}$— and a tropical basis $\mathcal{T}_{0}$ of $I_{\mathrm{proj}}$ consisting of polynomials of degree at most $d_{0}$. We can then take $\mathcal{T}$ to be the union of $\mathcal{G}$ with the set of polynomials from $\mathcal{T}_{0}$ evaluated in $x_{0}=1$; we will show that $\mathcal{T}$ is a tropical basis, from which it follows that we can take as $H$ the set of exponents in $\mathbb{N}^{n}$ of degree at most $d_{0}$.
Equip $K$ with a cross-section, so that we may define initial forms. By [MS15, Theorem 3.2.3], $\mathcal{T}$ is a tropical basis if and only if, for every $\gamma$, whenever there is a monomial in $\operatorname{in}_{\gamma}(I)$, then there is some $g\in\mathcal{T}$ such that $\operatorname{in}_{\gamma}(g)$ is a monomial.
Let $\gamma\in\Gamma^{n}$, and assume that $\operatorname{in}_{\gamma}(I)$ contains a monomial. Then by [MS15, Proposition 2.6.1], $\operatorname{in}_{0,\gamma}(I_{\mathrm{proj}})$ also contains a monomial, so there is $f\in\mathcal{T}_{0}$ such that $\operatorname{in}_{(0,\gamma)}(f)$ is a monomial. Since $\operatorname{in}_{(0,\gamma)}(f(x_{0},\ldots,x_{n}))=\operatorname{in}_{\gamma}(f(1,x_{1},\ldots,x_{n}))$ as shown in the proof of [MS15, Proposition 2.6.1], $\mathcal{T}$ is then a tropical basis. ∎
@@
This is [MS15, Lemma 2.6.2(1)]. One sees by inspection that in its proof (and in the proofs of Lemma 2.4.2 and Proposition 2.6.1, which are used within it) the rank $1$ assumption is never used. ∎
@@
    For all $f\in I$, the point $\gamma$ is a tropical root of $f$.
@@
By Proposition˜1.10, in $L$ we have that $\varphi_{H}(t,\gamma)$ holds if and only if $\gamma$ satisfies property (1) for $I_{t}$; by the Fundamental Theorem of Tropical Geometry [MS15, Theorem 3.2.3], in rank $1$ properties (1) and (3) are equivalent, so in $L$ we have that $\varphi_{H}(t,\gamma)$ holds if and only if $\gamma$ satisfies property (3) for $I_{t}$. This equivalence is a first-order statement, so it holds in algebraically closed nontrivially valued field of arbitrary rank by completeness of $\mathsf{ACVF}_{p,q}$. ∎
The following proposition, which generalises a step of the proof of the Fundamental Theorem of Tropical Geometry [MS15, Proposition 3.2.11], will be needed later on. We consider initial forms as computed with respect to a given angular component map, see ˜1.8.
@@
Let $V\subseteq\mathbb{G}_{\mathrm{m}}^{n}$ be an algebraic subvariety defined over an algebraically closed $\mathrm{ac}$-valued field $K$ with corresponding ideal $I\subseteq K[x^{\pm 1}]$, and $\gamma\in\Gamma^{n}$. Assume that $\operatorname{in}_{\gamma}(I)$ has a zero $\alpha$. Then the set of $a\in V(K)$ such that $v(a)=\gamma$ and $\operatorname{ac}(a)=\alpha$ is Zariski dense in $V(K)$.
@@
Without loss of generality the valuation is nontrivial. Fix a polynomial $g\in K[x^{\pm 1}]$. Let $H$ be given by Proposition˜1.10 applied to the set of exponents of a basis of $I$. As in the proof of ˜1.12, for every set of exponents $S$, we can write a first-order formula saying “for every variety $V$ defined by polynomials with exponents in $H$, every polynomial $g$ with exponents in $S$, every $\gamma\in\Gamma^{n}$, and every $\alpha\in\operatorname{in}_{\gamma}(V)$, if there are points $z_{0},z_{1}\in V$ with $v(z_{0})=\gamma$, $\operatorname{ac}(z_{0})=\alpha$ and $g(z_{1})\neq 0$, then there is a point $z_{2}\in V$ such that $v(z_{2})=\gamma$, $\operatorname{ac}(z_{2})=\alpha$, and $g(z_{2})\neq 0$”. This is first-order expressible and it holds in every algebraically closed $\mathrm{ac}$-valued field of rank 1 by [MS15, Proposition 3.2.11], so it holds in every algebraically closed nontrivially valued field by completeness of $\mathsf{ACVF}_{p,q,\mathrm{ac}}$. ∎
@@ 2 Purity and reducts
As already recalled, every algebraically closed valued field admits a cross-section of the valuation, see for example [ADH17, Lemma 3.3.32]. In this section, we show that universal domains for tropical difference algebra admit equivariant cross-sections of the valuation.
@@
Recall that a formula is existential iff it is of the form $\exists x\;\varphi(x)$ with $\varphi(x)$ quantifier-free, and positive (resp., positive primitive) iff it is obtained from atomic formulas by using $\land,\lor,\exists$ (resp., $\land,\exists$).
@@
    Let $\kappa$ be an infinite cardinal. A structure $M$ is existentially (resp., positively, resp., quantifier-free) $\kappa$-saturated iff, whenever $\pi(x)$ is a partial type consisting of existential (resp., positive, resp., quantifier-free) formulas with parameters from a set $A\subseteq M$ of size $\lvert A\rvert<\kappa$, there is $a\in M$ satisfying all formulas in $\pi(x)$.
@@
By standard model-theoretic facts, for every cardinal $\kappa$, every valued difference field has an existentially $\kappa$-saturated existentially closed extension.
@@
By a standard model-theoretic fact we may embed the valued field $(K,v)$ into a $\lvert K\rvert^{+}-$strongly homogeneous valued field $\mathfrak{U}$, that is, one where every elementary map with domain of size at most $\lvert K\rvert$ extends to an automorphism. Recall that, whenever a theory has quantifier elimination, an elementary map is nothing more than a partial isomorphism, and that quantifier elimination holds for $\mathsf{ACVF}$ in $\mathcal{L}_{\mathrm{vf}}$, see [HHM06, Theorem 2.1.1(iii)] [^4] Our language is slightly different, see 1.5..
@@ 2.1 Purity
Recall that, if $R$ is a ring, a submodule $N$ of an $R$-module $M$ is pure iff, for every matrix $A\in\operatorname{Mat}_{d\times n}(R)$ and vector $b\in N^{d}$, if the equation $A\cdot x=b$ has a solution in $M$ then it has a solution in $N$. This passes to localisations, and is in fact a local property, see [Lam99, §4, Exercise 34]. The case of interest for us is that purity as a $\mathbb{Z}[\sigma]$-module implies purity as a $\mathbb{Z}[\sigma,\sigma^{-1}]$-module, hence it suffices to work with $\mathbb{Z}[\sigma]$.
@@
Consider the $\mathbb{R}$-exponentiation (see [KS22, Definition 3.3.1] [^6] The reader familiar with this work will see that, if we decompose the automorphism group of $\mathbb{C}((\mathbb{R}))$ as in [KS22, Theorem 3.7.1] with respect to the natural cross-section, then any counterexample to purity must have a nontrivial factor in $\operatorname{Hom}(\mathbb{R},\mathbb{C}^{\times})$.) $(())\sigma\colon\mathbb{C}$$(\mathbb{R}$$)\to\mathbb{C}$$(\mathbb{R}$)) given by
@@
It suffices to let ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{n_{2}}$ be generators of the unique primitive lattice containing the one generated by ${\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{n_{1}}$ and of the same rank. See [Sch25, Lemma A.7]. ∎
@@
By [Sch25, Theorem A.13] and connectedness, $G$ is defined by a set of equations of the form $z^{\mathbf{u}}-1=0$ and the associated lattice $\Lambda\subseteq\mathbb{Z}^{n}$ is primitive. The subvariety $\operatorname{in}_{0}(t\cdot G)$ of $\mathbb{G}_{\mathrm{m}}^{n}(k)$ is contained in the coset by $\operatorname{res}(t)$ of the group $H$ defined by the equations $\operatorname{in}_{0}(z^{\mathbf{u}}-1)=0$, and it is easily checked that $\operatorname{in}_{0}(z^{\mathbf{u}}-1)=z^{\mathbf{u}}-1$. The lattice $\Lambda^{\prime}$ associated to $H$ must satisfy $\Lambda^{\prime}\supseteq\Lambda$. Moreover, $\operatorname{in}_{0}(t\cdot G)\subseteq\operatorname{res}(t)\cdot H$, and by [MS15, Lemma 3.2.6] [^7] This is for rank $1$ algebraically closed valued fields, but note that every algebraic group in this lemma is defined over the prime field., $\dim(\operatorname{in}_{0}(G))=\dim(G)$. It follows that $\dim(H)=\dim(G)$, hence $\Lambda^{\prime}$ and $\Lambda$ have the same rank, and because $\Lambda$ is primitive $\Lambda=\Lambda^{\prime}$. ∎
@@
Let $(K,\sigma)$ be an algebraically closed difference field, and let $I\subseteq K[y_{jh}^{\pm 1}\mid j\in[1,n],h\in[0,\ell]]$ be a proper ideal such that the variety $V(I)$ is a coset of an algebraic subgroup of $\mathbb{G}_{\mathrm{m}}^{n(\ell+1)}$. Assume moreover that there is $\alpha\in(K^{\times})^{n}$ such that, for every $r\leq s\in\mathbb{Z}$, we have that $\operatorname{orb}_{r,s}(\alpha)\in V(\sum_{p\in[r,s]}\sigma_{\to}^{p}(I))$.
@@
Finally, for (3) note that, under these assumptions, $I$ is generated by polynomials of the form $x^{\mathbf{u}}-b$, with $b\in\mathcal{O}^{\times}$. It follows that so is $I_{r,s}^{\prime}$. Different irreducible components of a subgroup of $\mathbb{G}_{\mathrm{m}}^{N}$ are translates of each other by a root of unity. As the latter have valuation 0, if one of them intersects $(\mathcal{O}^{\times})^{N}$, then so do all of them, and the conclusion follows. ∎
@@
    The formulas saying that $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is a $K$-generic point [^8] Recall that a point of an irreducible algebraic variety is $K$-generic iff it does not lie in any proper subvariety definable over $K$. of the irreducible algebraic variety $H_{r,s}$, all of whose coordinates have valuation 0.

### v2

@@ 1.2 Tropical geometry
In what follows, we will use results from tropical geometry, for which we refer the reader to [MS]. In the cited source, all value groups are Archimedean, which is typically not the case in this paper. Nevertheless, all the results we will use from [MS] still hold in our setting; since this sometimes requires justification, and because it may be of independent interest, we gather in this subsection the results that we will need. We especially want to highlight that many of these are, or can be reduced to, first-order statements; therefore, by the model-theoretic properties of $\mathsf{ACVF}$, they do not depend on the rank of the valuation.
@@
    The tropicalisation $\operatorname{trop}(f)\colon\Gamma^{n}\to\Gamma$ is defined as
@@
    The initial form of $f$ with respect to $\gamma$ is the Laurent polynomial over the residue field $k$ defined as $\operatorname{in}_{\gamma}(f)\coloneqq\sum_{\mathbf{u}\in S^{\prime}}\operatorname{res}(c_{\mathbf{u}}s(-v(c_{\mathbf{u}})))x^{\mathbf{u}}$, where $S^{\prime}\coloneqq\{\mathbf{u}\in S\mid\forall\mathbf{u}^{\prime}\in S,v(c_{\mathbf{u}})+\langle\mathbf{u},\gamma\rangle\leq v(c_{\mathbf{u}^{\prime}})+\langle\mathbf{u}^{\prime},\gamma\rangle\}$.
@@
    If $V\subseteq\mathbb{G}_{\mathrm{m}}^{n}(K)$ is an algebraic subvariety and $I$ is the ideal of Laurent polynomials over $K$ vanishing on $V$, then we denote by $\operatorname{in}_{\gamma}(V)$ the algebraic subvariety of $\mathbb{G}_{\mathrm{m}}^{n}(k)$ associated to $\operatorname{in}_{\gamma}(I)$. We call this the initial variety of $V$ with respect to $\gamma$.
Note that initial forms depend on the choice of a cross-section, even if this is suppressed from the notation. Below, if we mention initial forms without specifying a cross-section, we assume implicitly that a one exists and has been fixed. Recall that cross-sections always exist if the value group is divisible, e.g. in the algebraically closed case, see for example [aschenbrennerAsymptoticDifferentialAlgebra2017, Lemma 3.3.32].
@@
    In order to define initial forms one does not really need the cross-section $s$, rather the corresponding angular component $\operatorname{res}(xs(-v(x)))$. We mostly use the cross-section as that is customary in the literature, but observe that many results in tropical geometry (say from [MS, Sections 2 and 3]) generalise to initial forms when they are defined using an arbitrary angular component.
@@
It is almost immediate to see that the codimension 1 case of the Fundamental Theorem of Tropical Geometry, also known as Kapranov’s Theorem [MS, Theorem 3.1.3], is a first-order statement. In higher codimension we will need a bound of Joswig and Schröter on the minimum degree of a tropical basis.
@@
Let $K$ be a valued field.
  1. (1)
    Let $I$ be a homogeneous ideal of $K[{x}_{0},\ldots,{x}_{n}]$. A universal Gröbner basis $\mathcal{U}$ of $I$ is a finite subset of $I$ such that, for every $\gamma\in\Gamma(K)^{n+1}$, the set $\operatorname{in}_{\gamma}(\mathcal{U})$ generates $\operatorname{in}_{\gamma}(I)$.
  2. (2)
    Let $I$ be an ideal in $K[x^{\pm 1}]$. A tropical basis of $I$ is a finite generating subset $\mathcal{T}\subseteq I$ such that for every $\gamma\in\Gamma^{n}$, if there is $f\in I$ such that $\gamma$ is not a tropical root of $\operatorname{trop}(f)$, then there is $g\in\mathcal{T}$ such that $\gamma$ is not a tropical root of $\operatorname{trop}(g)$.
Given the ideal $I$, denote by $I_{\mathrm{proj}}$ its homogenisation: the homogeneous ideal in $K[x_{0},\ldots,x_{n}]$ obtained by homogenising the elements of $K[x^{\pm 1}]\cap I$.
@@
Let $K$ be an algebraically closed valued field with value group $\mathbb{Q}$. For all finite $S\subseteq\mathbb{Z}^{n}$ there are $d\in\mathbb{N}$ and a finite $H\subseteq\mathbb{N}^{n}$ such that the following holds. Let $I\subseteq K[x^{\pm 1}]$ be an ideal and $\mathcal{G}$ a finite set generating of $I$ such that every polynomial in $\mathcal{G}$ can be written as $\sum_{\mathbf{u}\in S}c_{\mathbf{u}}x^{\mathbf{u}}$ (with possibly some $c_{\mathbf{u}}=0$).
  1. (1)
    There is a universal Gröbner basis $\mathcal{U}$ of $I_{\mathrm{proj}}$ such that every polynomial in $\mathcal{U}$ can be written as $\sum_{\mathbf{u}\in\{0,\ldots,d\}\times H}b_{\mathbf{u}}x^{\mathbf{u}}$.
  2. (2)
    There is a tropical basis $\mathcal{T}\supseteq\mathcal{G}$ of $I$ such that every polynomial in $\mathcal{T}$ can be written as $\sum_{\mathbf{u}\in H}b_{\mathbf{u}}x^{\mathbf{u}}$.
@@
Because $\Gamma(K)=\mathbb{Q}$, every ideal of $K[x_{0},\ldots,x_{n}]$ is defined over a subfield of $K$ with discrete valuation. Hence we may apply [joswigDegreeTropicalBasis2018, Theorem 10] to $I_{\mathrm{proj}}$ and obtain that there are $d$, depending only on $n$ and on the minimal degree of a set of generators of $I_{\mathrm{proj}}$, a tropical basis $\mathcal{T}_{0}$, and a universal Gröbner basis $\mathcal{U}$ of $I_{\mathrm{proj}}$, both consisting of polynomials of degree at most $d$. We can then take $\mathcal{T}$ to be the union of $\mathcal{G}$ with the set of polynomials from $\mathcal{T}_{0}$ evaluated in $x_{0}=1$; we will show that $\mathcal{T}$ is a tropical basis, from which it follows that we can take as $H$ the set of exponents $0,d]^{n}\subseteq\mathbb{N}^{n}$.
Equip $K$ with a cross-section, so that we may define initial forms. By [MS, Theorem 3.2.3], $\mathcal{T}$ is a tropical basis if and only if, for every $\gamma$, whenever there is a monomial in $\operatorname{in}_{\gamma}(I)$, then there is some $g\in\mathcal{T}$ such that $\operatorname{in}_{\gamma}(g)$ is a monomial.
Let $\gamma\in\Gamma^{n}$, and assume that $\operatorname{in}_{\gamma}(I)$ contains a monomial. Then by [MS, Proposition 2.6.1], $\operatorname{in}_{(0,\gamma)}(I_{\mathrm{proj}})$ also contains a monomial, so there is $f\in\mathcal{T}_{0}$ such that $\operatorname{in}_{(0,\gamma)}(f)$ is a monomial. Since $\operatorname{in}_{(0,\gamma)}(f(x_{0},\ldots,x_{n}))=\operatorname{in}_{\gamma}(f(1,x_{1},\ldots,x_{n}))$ as shown in the proof of [MS, Proposition 2.6.1], $\mathcal{T}$ is then a tropical basis. ∎
@@
This is [MS, Lemma 2.6.2(1)]. One sees by inspection that in its proof (and in the proofs of Lemma 2.4.2 and Proposition 2.6.1, which are used within it) the rank $1$ assumption is never used. ∎
@@
    For all $f\in I$, the point $\gamma$ is a tropical root of $\operatorname{trop}(f)$.
@@
By Proposition˜1.10, in $L$ we have that $\varphi_{H}(t,\gamma)$ holds if and only if $\gamma$ satisfies property (1) for $I_{t}$; by the Fundamental Theorem of Tropical Geometry [MS, Theorem 3.2.3], in rank $1$ properties (1) and (3) are equivalent, so in $L$ we have that $\varphi_{H}(t,\gamma)$ holds if and only if $\gamma$ satisfies property (3) for $I_{t}$. This equivalence is a first-order statement, so it holds in algebraically closed nontrivially valued fields of arbitrary rank by completeness of $\mathsf{ACVF}_{p,q}$. ∎
The following proposition, which generalises a step of the proof of the Fundamental Theorem of Tropical Geometry [MS, Proposition 3.2.11], will be needed later on. We consider initial forms as computed with respect to a given angular component map, see ˜1.8.
@@
Let $V\subseteq\mathbb{G}_{\mathrm{m}}^{n}$ be an algebraic subvariety defined over an algebraically closed nontrivially $\mathrm{ac}$-valued field $K$, and $\gamma\in\Gamma^{n}$.
  1. (1)
    The variety $\operatorname{in}_{\gamma}(V)(k(K))$ coincides with the set of angular components of points of $V(K)$ of valuation $\gamma$.
  2. (2)
    For each $\alpha\in\operatorname{in}_{\gamma}(V)(k(K))$, the set of $a\in V(K)$ such that $v(a)=\gamma$ and $\operatorname{ac}(a)=\alpha$ is Zariski dense in $V(K)$.
@@
Let $I\subseteq K[x^{\pm 1}]$ be the ideal associated to $V$. Let $d$ and $H$ be given by Proposition˜1.10 applied to the set $S$ of exponents of a fixed basis of $I$. Similarly as in the proof of ˜1.12, write $I=I_{t}$ where $t$ is the tuple of coefficients of a set of generators with exponents in $S$. Let $J_{t,\gamma}$ be the ideal generated by $\operatorname{in}_{\gamma}(f)$, as $f$ varies amongst polynomials in $I_{t}$ with exponents in $H$. In every nontrivially $\operatorname{ac}$-valued field, for every $t$ and $\gamma$, it follows from the definition of initial form that
$\{\operatorname{ac}(a)\mid a\in V(I_{t}),v(a)=\gamma\}\subseteq V(\operatorname{in}_{\gamma}(I_{t}))\subseteq V(J_{t,\gamma}).$
By [MS, Proposition 3.2.11], in every algebraically closed $\mathrm{ac}$-valued field of rank 1, the first inclusion above is an equality. If furthermore $\Gamma(K)=\mathbb{Q}$, then by Proposition˜1.10 there is a universal Gröbner basis of $I_{\mathrm{proj}}$ with exponents in $\{0,\ldots,d\}\times H$, and it follows that the second inclusion is also an equality.
As the sets $\{\operatorname{ac}(a)\mid a\in V(I_{t}),v(a)=\gamma\}$ and $V(J_{t,\gamma})$ are first-order definable uniformly in $t$ and $\gamma$, completeness of $\mathsf{ACVF}_{p,q,\mathrm{ac}}$ yields that they coincide in every algebraically closed nontrivially $\operatorname{ac}$-valued field. As $V(\operatorname{in}_{\gamma}(I_{t}))$ lies between them, this proves (1).
As for (2), fix a polynomial $g\in K[x^{\pm 1}]$. The conclusion follows from the previous point and the first-order expressibility of “if there are points $z_{0},z_{1}\in V$ with $v(z_{0})=\gamma$, $\operatorname{ac}(z_{0})=\alpha$ and $g(z_{1})\neq 0$, then there is a point $z_{2}\in V$ such that $v(z_{2})=\gamma$, $\operatorname{ac}(z_{2})=\alpha$, and $g(z_{2})\neq 0$”. ∎
@@ 2 Purity and reducts
As already recalled, every algebraically closed valued field admits a cross-section of the valuation. In this section, we show that universal domains for tropical difference algebra admit one that is $\sigma$-equivariant.
@@
Recall that a formula is existential iff it is of the form $\exists x\;\varphi(x)$ with $\varphi(x)$ quantifier-free, and positive (resp. positive primitive) iff it is obtained from atomic formulas by using $\land,\lor,\exists$ (resp. $\land,\exists$).
@@
    Let $\kappa$ be an infinite cardinal. A structure $M$ is existentially (resp. positively, resp. quantifier-free) $\kappa$-saturated iff, whenever $\pi(x)$ is a partial type consisting of existential (resp. positive, resp. quantifier-free) formulas with parameters from a set $A\subseteq M$ of size $\lvert A\rvert<\kappa$, there is $a\in M$ satisfying all formulas in $\pi(x)$.
@@
By standard model-theoretic arguments, for every cardinal $\kappa$, every valued difference field has an existentially $\kappa$-saturated existentially closed extension.
@@
As commonly done in model theory, we embed the valued field $(K,v)$ into a $\lvert K\rvert^{+}-$strongly homogeneous valued field $\mathfrak{U}$, that is, one where every elementary map with domain of size at most $\lvert K\rvert$ extends to an automorphism. Recall that, whenever a theory has quantifier elimination, an elementary map is nothing more than a partial isomorphism, and that quantifier elimination holds for $\mathsf{ACVF}$ in $\mathcal{L}_{\mathrm{vf}}$, see [haskellDefinableSetsAlgebraically2006, Theorem 2.1.1(iii)] [^3] Our language is slightly different, see 1.5..
@@ 2.1 Purity
Recall that, if $R$ is a ring, a submodule $N$ of an $R$-module $M$ is pure iff, for every matrix $A\in\operatorname{Mat}_{d\times n}(R)$ and vector $b\in N^{d}$, if the equation $A\cdot x=b$ has a solution in $M$ then it has a solution in $N$. This passes to localisations, and is in fact a local property, see [lamLecturesModulesRings1999, §4, Exercise 34]. The case of interest for us is that purity as a $\mathbb{Z}[\sigma]$-module implies purity as a $\mathbb{Z}[\sigma,\sigma^{-1}]$-module, hence it suffices to work with $\mathbb{Z}[\sigma]$.
@@
Consider the $\mathbb{R}$-exponentiation (see [kuhlmannAutomorphismGroupValued2022, Definition 3.3.1] [^5] The reader familiar with this work will see that, if we decompose the automorphism group of $\mathbb{C}((\mathbb{R}))$ as in [kuhlmannAutomorphismGroupValued2022, Theorem 3.7.1] with respect to the natural cross-section, then any counterexample to purity must have a nontrivial factor in $\operatorname{Hom}(\mathbb{R},\mathbb{C}^{\times})$.) $(())\sigma\colon\mathbb{C}$$(\mathbb{R}$$)\to\mathbb{C}$$(\mathbb{R}$)) given by
@@
It suffices to let ${\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{n_{2}}$ be generators of the unique primitive lattice containing the one generated by ${\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{n_{1}}$ and of the same rank. See [scheferCountingTorsionPoints2025, Lemma A.7]. ∎
@@
By [scheferCountingTorsionPoints2025, Theorem A.13] and connectedness, $G$ is defined by a set of equations of the form $z^{\mathbf{u}}-1=0$ and the associated lattice $\Lambda\subseteq\mathbb{Z}^{n}$ is primitive. The subvariety $\operatorname{in}_{0}(t\cdot G)$ of $\mathbb{G}_{\mathrm{m}}^{n}(k)$ is contained in the coset by $\operatorname{res}(t)$ of the group $H$ defined by the equations $\operatorname{in}_{0}(z^{\mathbf{u}}-1)=0$, and it is easily checked that $\operatorname{in}_{0}(z^{\mathbf{u}}-1)=z^{\mathbf{u}}-1$. The lattice $\Lambda^{\prime}$ associated to $H$ must satisfy $\Lambda^{\prime}\supseteq\Lambda$. Moreover, $\operatorname{in}_{0}(t\cdot G)\subseteq\operatorname{res}(t)\cdot H$, and by [MS, Lemma 3.2.6] [^6] This is for rank $1$ algebraically closed valued fields, but note that every algebraic group in this lemma is defined over the prime field., $\dim(\operatorname{in}_{0}(G))=\dim(G)$. It follows that $\dim(H)=\dim(G)$, hence $\Lambda^{\prime}$ and $\Lambda$ have the same rank, and because $\Lambda$ is primitive $\Lambda=\Lambda^{\prime}$. ∎
@@
Let $(K,\sigma)$ be an algebraically closed difference field, and let $I\subseteq K[y_{jh}^{\pm 1}\mid j\in[1,n],h\in[0,\ell]]$ be a proper ideal such that the variety $V(I)$ is a coset of an algebraic subgroup of $\mathbb{G}_{\mathrm{m}}^{n(\ell+1)}$. Assume moreover that there is $\alpha\in(K^{\times})^{n}$ such that, for all $r\leq s\in\mathbb{Z}$, we have that $\operatorname{orb}_{r,s}(\alpha)\in V(\sum_{p\in[r,s]}\sigma_{\to}^{p}(I))$.
@@
Finally, for (3) note that, under these assumptions, $I$ is generated by polynomials of the form $x^{\mathbf{u}}-b$, with $b\in\mathcal{O}^{\times}$. It follows that so is $I_{r,s}^{\prime}$, so $V(I_{r,s}^{\prime})$ has a point with all coordinates in $\mathcal{O}^{\times}$ by ˜2.5. Different irreducible components of a subgroup of $\mathbb{G}_{\mathrm{m}}^{N}$ are translates of each other by a root of unity. As roots of unity have valuation 0, if one of them intersects $(\mathcal{O}^{\times})^{N}$, then so do all of them, and the conclusion follows. ∎
@@
    The formulas saying that $(y_{jh})_{j\in[1,n],h\in[r,s+\ell]}$ is a $K$-generic point [^7] Recall that a point of an irreducible algebraic variety is $K$-generic iff it does not lie in any proper subvariety definable over $K$. of the irreducible variety $H_{r,s}$, all of whose coordinates have valuation 0.